The ratio of the number of boys and girls in a school is 3:2 Out of these 10% of the boys and 25% of girls are scholarship holders. % of students who are not scholarship holders.?
Answer by richwmiller(17219)

there are 5x students
3x are boys and 2x are girls
3x/10 have scholarships and 2x/4 have scholarship
3x/10+5x/10=8x/10 have scholarships
5x-.8x=4.2x have no scholarships
4.2/5=8.4/10=.84 =84% do not have scholarships

or .8x/5x=.16 = 16% have scholarships 100%-16%=84% have no scholarships